# TurkBench: A Benchmark for Evaluating Turkish Large Language Models

### v1

@@ 1 Introduction
Despite the success of these evaluation frameworks, a significant gap exists in their linguistic and cultural diversity. The majority of such benchmarks are predominantly English-centric, which stems from the availability of high-quality digital text in English Lai et al. (2024). This creates a critical challenge for evaluating models in other languages Park et al. (2024); Nacar et al. (2025); Baucells et al. (2025); Magnini et al. (2025), as direct translation of English benchmarks often fails to capture the unique linguistic structures, cultural nuances, and specific knowledge inherent to other languages Umutlu et al. (2025). Turkish, with its agglutinative morphology and distinct syntactic rules, presents unique challenges that cannot be easily assessed by frameworks designed for Indo-European languages Hakkani-Tür et al. (2002); Oflazer (2014).
Although there are efforts to develop benchmarks for Turkish LLMs Safaya et al. (2022); Uzunoğlu and Şahin (2023); Yüksel et al. (2024); Er et al. (2025), these resources primarily focus on exam-style tasks or do not cover many capabilities that drive real-world use of Turkish LLMs, such as open-ended instruction following, safety and content moderation, or fine-grained grammar and vocabulary control. Moreover, such benchmarks are mostly designed by using existing publicly available datasets, and lack of public leaderboards serving automated evaluation. In this study, we present TurkBench, which is a comprehensive benchmark specifically designed by curating novel data for the Turkish language. TurkBench provides a robust, culturally-aware, and linguistically-sound online evaluation tool to accurately measure the true capabilities of LLMs in Turkish.
@@ 2 Related Work
Mukayese is one of the first benchmarks for Turkish, which assembles datasets for tasks such as text classification, named entity recognition, and sentiment analysis Safaya et al. (2022). TurkishMMLU Yüksel et al. (2024) and TR-MMLU Bayram et al. (2025) adapt the MMLU paradigm to Turkish language, providing large-scale multiple-choice question answering benchmarks that reflect national educational standards; yet these resources primarily focus on exam-style multiple-choice question tasks. TR-MTEB focuses on sentence embeddings and introduces the first large-scale benchmark for Turkish sentence representations, covering six task families over 26 datasets Baysan et al. (2025). Turkish-PLU (Procedural Language Understanding) constructs a corpus of Turkish WikiHow texts, and defines tasks such as action linking and summarization; showing that Turkish-specific models outperform multilingual baselines on procedural understanding tasks Uzunoğlu and Şahin (2023). Cetvel is a recent benchmark covering 23 tasks grouped into seven categories based on existing publicly available datasets Er et al. (2025). Compared to these efforts, TurkBench differs in several key aspects. (i) All 8,151 instances across 21 subtasks are newly curated from high-quality Turkish sources with strategic partnerships with national institutions and university departments. (ii) All data is validated by human experts for correctness, grammatical well-formedness, and cultural sensitivity. (iii) We cover not only standard language understanding and classification but also reasoning, safety-oriented content moderation, and instruction following, which provides a much broader view of Turkish LLM capabilities. (iv) TurkBench is integrated into a public leaderboard that supports near-automatic submission and standardized evaluation.
@@ 4.1 Experimental Design
Our benchmark covers a total of 21 tasks, consisting of 7 open-ended and 14 multiple-choice tasks. For the open-ended tasks, we adopt an LLM-as-a-Judge paradigm. In particular, we employ GPT-4o-mini as the judge model, which evaluates system responses based on semantic alignment with reference answers. This approach enables us to assess inherently subjective tasks (e.g., summarization, faithfulness, instruction following) that cannot be reduced to a single ground truth label. For the multiple-choice tasks, in contrast, objective evaluation is performed by directly comparing model predictions against gold answers.
@@ 4.3 Experimental Results
Model TK ML RC NL SM ST MA CX CS BI TR FA TX IF SA TD VO NE PS ID Avg
gpt-oss-120b 77.6 72.2 96.1 80.0 81.8 90.4 26.6 80.8 94.7 82.8 75.3 85.1 99.9 93.6 48.9 85.6 85 70.0 78.2 68.0 78.6
GLM-4.6 83.9 76.0 94.0 85.5 76.5 88.2 34.4 48.1 63.1 88.6 67.0 87.8 95.6 92.6 49.6 84.4 96 69.2 79.0 78.0 76.9
DeepSeek-V3.1 77.6 46.7 95.1 28.2 80.1 91.5 58.6 66.3 91.8 79.6 77.2 87.4 98.7 94.9 48.9 84.8 93 56.5 83.6 63.3 75.2
Qwen3-Next-80B-Inst 67.2 61.4 94.4 85.1 81.7 93.0 63.4 26.9 92.2 74.2 67.9 88.7 95.3 94.3 46.7 85.6 80 58.3 81.3 61.3 75.0
Qwen3-30B-A3B-Inst 65.1 58.9 90.8 81.6 81.9 86.9 55.4 22.1 90.6 78.2 62.9 83.4 97.5 92.5 44.5 95.2 89 55.0 78.1 68.7 73.4
gemma-3-27b-it 74.7 53.5 93.0 75.7 74.1 90.2 19.6 27.9 91.4 80.1 61.8 84.4 94.0 84.8 42.6 88.1 90 66.0 84.4 85.3 73.0
Qwen3-235B-Instruct 73.4 61.6 95.4 33.7 81.8 92.2 47.2 24.0 91.4 73.2 77.4 89.1 98.1 94.5 47.4 87.7 77 58.5 81.0 59.3 72.2
gemma-3-12b-TR-V1 71.4 48.0 92.6 71.4 77.0 92.6 23.4 26.9 91.0 77.8 60.1 86.2 93.0 83.7 46.7 80.3 90 63.9 73.4 74.7 71.2
gemma-3-12b-it 71.4 48.2 92.6 71.4 77.5 92.6 22.4 23.1 91.4 78.0 59.6 88.1 92.7 81.2 47.5 79.8 91 63.5 73.6 74.7 71.0
Qwen3-235B-A22B 66.1 61.3 94.0 29.4 71.5 93.3 33.8 15.4 88.5 76.0 60.0 86.1 98.7 91.2 43.1 86.4 57 55.8 77.7 79.3 68.2
Qwen2.5-14B-Instruct 58.9 45.2 89.0 82.4 75.0 91.9 18.8 27.9 90.2 82.8 51.1 83.4 99.0 64.5 46.0 68.7 80 51.6 64.3 60.0 66.5
Tongyi-DR-30B 64.6 35.3 85.6 75.7 72.9 88.0 29.0 1.0 85.7 84.2 59.5 85.6 98.3 59.9 46.0 76.1 83 55.1 54.0 68.0 65.4
TR-Gemma-9b-v0.1 67.7 48.0 92.5 62.0 78.3 89.1 9.8 23.1 89.3 79.2 65.1 88.3 96.7 83.3 46.0 80.3 84 0.0 46.8 77.3 65.3
Qwen3-32B 66.7 38.0 93.0 81.6 72.8 90.9 0.0 26.0 91.0 8.6 61.0 87.6 98.0 83.8 48.9 85.2 71 60.4 77.7 50.7 64.6
gemma-2-9b-it 62.3 39.4 90.3 76.5 74.4 86.8 6.4 8.7 89.3 18.8 54.2 86.0 96.1 73.6 39.5 84.8 81 48.8 69.1 68.0 62.7
aya-expanse-8b 55.7 32.5 90.0 40.8 77.0 81.1 4.2 23.1 86.9 24.2 54.8 87.0 98.4 82.2 41.6 71.6 78 38.6 63.9 40.0 58.6
Qwen2.5-7B-Instruct 39.1 29.4 85.3 53.7 73.4 91.8 10.8 23.1 74.6 68.2 43.4 82.4 97.9 61.8 23.4 70.0 39 34.8 63.8 32.7 54.9
Llama-3.1-8B-Instruct 40.1 19.0 89.5 31.0 71.5 81.9 2.8 25.0 33.2 36.8 45.3 79.2 98.6 71.9 21.2 28.4 38 26.7 61.6 13.3 45.7
DeepSeek-R1-Q3-8B 31.3 20.9 84.0 32.2 68.3 54.1 0.2 21.2 50.8 12.0 55.7 80.0 98.5 73.1 22.6 39.5 24 44.5 45.8 22.0 44.0
Qwen3-14B 35.4 29.5 72.2 32.2 68.3 61.2 14.6 5.8 80.7 44.0 21.1 66.7 98.0 32.1 30.7 58.8 44 19.7 18.0 26.7 43.0
Phi-4-mini-instruct 3.7 10.3 84.9 60.0 65.4 89.3 6.2 15.4 17.6 12.2 31.8 74.9 98.4 62.1 37.2 46.9 41 32.3 31.3 21.3 42.1
gemma-2-2b-it 24.5 6.4 86.1 32.2 70.2 84.2 2.0 18.3 76.2 33.2 37.6 83.4 96.6 59.1 0.0 0.4 0 31.9 29.4 1.3 38.6
Magistral-Small-2506 49.5 27.4 57.8 3.9 53.0 81.7 11.4 7.7 38.1 25.3 16.6 82.6 96.1 10.7 21.9 68.7 15 27.4 57.7 12.7 38.3
Qwen3-1.7B 27.4 21.3 52.5 41.2 62.6 51.0 0.1 20.2 55.3 10.9 22.3 82.7 98.1 10.1 30.9 17.3 23 36.6 40.0 26.0 36.5
TDM-8b-v0.1 23.4 21.3 51.3 34.1 66.8 0.0 2.0 18.3 58.2 0.0 25.2 87.8 95.7 5.0 24.1 29 0.0 0.0 27.3 30.0
Qwen3-0.6B 23.3 19.1 35.5 30.2 32.4 49.7 0.0 18.3 65.6 8.9 12.6 69.2 100.0 4.5 19.1 10.7 26 20.9 17.2 26.0 29.5
Kumru-2B 2.7 10.6 79.0 1.6 54.7 58.0 0.2 0.0 8.6 29.2 49.3 76.5 97.4 47.0 11.0 6.2 12 0.0 0.4 2.0 27.3
@@
Task Abbreviations: TK = Turkish General Knowledge, ML = MMLU Topics, RC = Reading Comprehension, NL = Natural Language Inference, SM = Summarization, ST = Semantic Textual Similarity, MA = Mathematics, CX = Complex Reasoning, CS = Commonsense Reasoning, BI = Bias (MC), TR = Truthfulness, FA = Faithfulness, TX = Toxicity, IF = Instruction Following, SA = Sentiment Analysis, TD = Topic Detection, VO = Turkish Vocabulary, NE = Named Entity Recognition, PS = Parts of Speech, ID = Metaphors & Idioms, Avg = Overall Average. Models are sorted by average score (descending).
@@
We evaluate and compare 27 open-source large language model in the TurkBench benchmark, sorted by the average score across all tasks, in Table 5. The model details are given in Appendix 9.4.
@@ Acknowledgments:
We would like to thank to OSYM (The Turkish Measuring, Selection and Placement Center), TUBITAK (The Scientific and Technological Research Council of Turkey), and Turkcell AI for their permissions to use certain resources in this benchmark.

### v2

@@ 1 Introduction
Despite the success of these evaluation frameworks, a significant gap exists in their linguistic and cultural diversity. The majority of such benchmarks are predominantly English-centric, which stems from the availability of high-quality digital text in English Lai et al. (2024). This creates a critical challenge for evaluating models in other languages Park et al. (2024); Nacar et al. (2025); Baucells et al. (2025); Magnini et al. (2025), as direct translation of English benchmarks often fails to capture the unique linguistic structures, cultural nuances, and specific knowledge inherent to other languages Umutlu et al. (2025). Turkish, with its agglutinative morphology and distinct syntactic rules, presents unique challenges that cannot be easily assessed by frameworks designed for Indo-European languages Hakkani-Tür et al. (2002); Oflazer (2014); Toraman (2024).
Although there are efforts to develop benchmarks for Turkish LLMs Safaya et al. (2022); Uzunoğlu and Şahin (2023); Yüksel et al. (2024); Er et al. (2025), these resources are mostly designed by using existing publicly available datasets, and lack of public leaderboards accepting online submissions for automated evaluation. Moreover, existing benchmarks do not cover a wide range of capabilities that drive real-world use of Turkish LLMs, such as open-ended instruction following, safety and content moderation, or fine-grained grammar and vocabulary control. In this study, we present TurkBench, which is a comprehensive benchmark specifically designed by curating novel data for the Turkish language. TurkBench provides a robust, culturally-aware, and linguistically-sound online evaluation tool to accurately measure the true capabilities of LLMs in Turkish.
@@ 2 Related Work
Mukayese is one of the first benchmarks for Turkish, which assembles datasets for tasks such as text classification, named entity recognition, and sentiment analysis Safaya et al. (2022). TurkishMMLU Yüksel et al. (2024) and TR-MMLU Bayram et al. (2025) adapt the MMLU paradigm to Turkish language, providing large-scale multiple-choice question answering benchmarks that reflect national educational standards; yet these resources primarily focus on exam-style multiple-choice question tasks. TR-MTEB focuses on sentence embeddings and introduces the first large-scale benchmark for Turkish sentence representations, covering six task families over 26 datasets Baysan et al. (2025). Turkish-PLU (Procedural Language Understanding) constructs a corpus of Turkish WikiHow texts, and defines tasks such as action linking and summarization; showing that Turkish-specific models outperform multilingual baselines on procedural understanding tasks Uzunoğlu and Şahin (2023). There are also some efforts to develop the Turkish versions of some existing English benchmarks Acikgoz et al. (2024). Cetvel is a recent benchmark covering 23 tasks grouped into seven categories based on existing publicly available datasets Er et al. (2025). Compared to these efforts, TurkBench differs in several key aspects. (i) All 8,151 instances across 21 subtasks are newly curated from high-quality Turkish sources with strategic partnerships with national institutions and university departments. (ii) All data is validated by human experts for correctness, grammatical well-formedness, and cultural sensitivity. (iii) We cover not only standard language understanding and classification but also reasoning, safety-oriented content moderation, and instruction following, which provides a much broader view of Turkish LLM capabilities. (iv) TurkBench is integrated into a public leaderboard that supports near-automatic submission and standardized evaluation.
@@ 4.1 Experimental Design
Our benchmark covers a total of 21 tasks, consisting of 7 open-ended and 14 multiple-choice tasks. For the open-ended tasks, we adopt an LLM-as-a-Judge paradigm. In particular, we employ GPT-4o-mini as the judge model, which evaluates system responses based on semantic alignment with reference answers. This approach enables us to assess inherently subjective tasks (e.g., summarization, faithfulness, instruction following) that cannot be reduced to a single ground truth label. For the multiple-choice tasks, in contrast, objective evaluation is performed by directly comparing model predictions against gold answers. We do not share the prompts used for LLM-as-a-Judge due to online leaderboard evaluation.
@@ 4.3 Experimental Results
Model Knowl. Lang. Understand. Reasoning Content Moderation Grammar IF Avg
TK ML RC NL SM ST MA CX CS BI TR FA TX SA TD VO NE PS ID IF
gpt-oss-120b 77.6 72.2 96.1 80.0 81.8 90.4 26.6 80.8 94.7 82.8 75.3 85.1 99.9 48.9 85.6 85 70.0 78.2 68.0 78.6 93.6
GLM-4.6 83.9 76.0 94.0 85.5 76.5 88.2 34.4 48.1 63.1 88.6 67.0 87.8 95.6 49.6 84.4 96 69.2 79.0 78.0 76.9 92.6
DeepSeek-V3.1 77.6 46.7 95.1 28.2 80.1 91.5 58.6 66.3 91.8 79.6 77.2 87.4 98.7 48.9 84.8 93 56.5 83.6 63.3 75.2 94.9
Qwen3-80B-Inst 67.2 61.4 94.4 85.1 81.7 93.0 63.4 26.9 92.2 74.2 67.9 88.7 95.3 46.7 85.6 80 58.3 81.3 61.3 75.0 94.3
Qwen3-30B-Inst 65.1 58.9 90.8 81.6 81.9 86.9 55.4 22.1 90.6 78.2 62.9 83.4 97.5 44.5 95.2 89 55.0 78.1 68.7 73.4 92.5
gemma-3-27b-it 74.7 53.5 93.0 75.7 74.1 90.2 19.6 27.9 91.4 80.1 61.8 84.4 94.0 42.6 88.1 90 66.0 84.4 85.3 73.0 84.8
Qwen3-235B-Inst 73.4 61.6 95.4 33.7 81.8 92.2 47.2 24.0 91.4 73.2 77.4 89.1 98.1 47.4 87.7 77 58.5 81.0 59.3 72.2 94.5
gemma-3-12b-TR 71.4 48.0 92.6 71.4 77.0 92.6 23.4 26.9 91.0 77.8 60.1 86.2 93.0 46.7 80.3 90 63.9 73.4 74.7 71.2 83.7
gemma-3-12b-it 71.4 48.2 92.6 71.4 77.5 92.6 22.4 23.1 91.4 78.0 59.6 88.1 92.7 47.5 79.8 91 63.5 73.6 74.7 71.0 81.2
Qwen3-235B 66.1 61.3 94.0 29.4 71.5 93.3 33.8 15.4 88.5 76.0 60.0 86.1 98.7 43.1 86.4 57 55.8 77.7 79.3 68.2 91.2
Qwen2.5-14B-Inst 58.9 45.2 89.0 82.4 75.0 91.9 18.8 27.9 90.2 82.8 51.1 83.4 99.0 46.0 68.7 80 51.6 64.3 60.0 66.5 64.5
Tongyi-DR-30B 64.6 35.3 85.6 75.7 72.9 88.0 29.0 1.0 85.7 84.2 59.5 85.6 98.3 46.0 76.1 83 55.1 54.0 68.0 65.4 59.9
TR-Gemma-9b 67.7 48.0 92.5 62.0 78.3 89.1 9.8 23.1 89.3 79.2 65.1 88.3 96.7 46.0 80.3 84 0.0 46.8 77.3 65.3 83.3
Qwen3-32B 66.7 38.0 93.0 81.6 72.8 90.9 0.0 26.0 91.0 8.6 61.0 87.6 98.0 48.9 85.2 71 60.4 77.7 50.7 64.6 83.8
gemma-2-9b-it 62.3 39.4 90.3 76.5 74.4 86.8 6.4 8.7 89.3 18.8 54.2 86.0 96.1 39.5 84.8 81 48.8 69.1 68.0 62.7 73.6
aya-expanse-8b 55.7 32.5 90.0 40.8 77.0 81.1 4.2 23.1 86.9 24.2 54.8 87.0 98.4 41.6 71.6 78 38.6 63.9 40.0 58.6 82.2
Qwen2.5-7B-Inst 39.1 29.4 85.3 53.7 73.4 91.8 10.8 23.1 74.6 68.2 43.4 82.4 97.9 23.4 70.0 39 34.8 63.8 32.7 54.9 61.8
Llama-3.1-8B-Inst 40.1 19.0 89.5 31.0 71.5 81.9 2.8 25.0 33.2 36.8 45.3 79.2 98.6 21.2 28.4 38 26.7 61.6 13.3 45.7 71.9
DeepSeek-Q3-8B 31.3 20.9 84.0 32.2 68.3 54.1 0.2 21.2 50.8 12.0 55.7 80.0 98.5 22.6 39.5 24 44.5 45.8 22.0 44.0 73.1
Qwen3-14B 35.4 29.5 72.2 32.2 68.3 61.2 14.6 5.8 80.7 44.0 21.1 66.7 98.0 30.7 58.8 44 19.7 18.0 26.7 43.0 32.1
Phi-4-mini-instruct 3.7 10.3 84.9 60.0 65.4 89.3 6.2 15.4 17.6 12.2 31.8 74.9 98.4 37.2 46.9 41 32.3 31.3 21.3 42.1 62.1
gemma-2-2b-it 24.5 6.4 86.1 32.2 70.2 84.2 2.0 18.3 76.2 33.2 37.6 83.4 96.6 0.0 0.4 0 31.9 29.4 1.3 38.6 59.1
Magistral-Small 49.5 27.4 57.8 3.9 53.0 81.7 11.4 7.7 38.1 25.3 16.6 82.6 96.1 21.9 68.7 15 27.4 57.7 12.7 38.3 10.7
Qwen3-1.7B 27.4 21.3 52.5 41.2 62.6 51.0 0.1 20.2 55.3 10.9 22.3 82.7 98.1 30.9 17.3 23 36.6 40.0 26.0 36.5 10.1
TDM-8b-v0.1 23.4 21.3 51.3 34.1 66.8 0.0 2.0 18.3 58.2 0.0 25.2 87.8 95.7 24.1 29 0.0 0.0 27.3 30.0 5.0
Qwen3-0.6B 23.3 19.1 35.5 30.2 32.4 49.7 0.0 18.3 65.6 8.9 12.6 69.2 100 19.1 10.7 26 20.9 17.2 26.0 29.5 4.5
Kumru-2B 2.7 10.6 79.0 1.6 54.7 58.0 0.2 0.0 8.6 29.2 49.3 76.5 97.4 11.0 6.2 12 0.0 0.4 2.0 27.3 47.0
@@
Task Abbreviations: TK = Turkish General Knowledge, ML = MMLU Topics, RC = Reading Comprehension, NL = Natural Language Inference, SM = Summarization, ST = Semantic Textual Similarity, MA = Mathematics, CX = Complex Reasoning, CS = Commonsense Reasoning, BI = Bias (MC), TR = Truthfulness, FA = Faithfulness, TX = Toxicity, SA = Sentiment Analysis, TD = Topic Detection, VO = Turkish Vocabulary, NE = Named Entity Recognition, PS = Parts of Speech, ID = Metaphors & Idioms, IF = Instruction Following, Avg = Overall Average. Models are sorted by average score (descending).
@@
We evaluate and compare 27 open-source large language model in the TurkBench benchmark, sorted by the average score across all tasks, in Table 5. The results are based on a single run of each model. The model details are given in Appendix 9.4.
@@ Acknowledgments:
We would like to thank to OSYM (The Turkish Measuring, Selection and Placement Center), TUBITAK (The Scientific and Technological Research Council of Turkey), and Turkcell AI for their permissions to use certain resources in this benchmark. We would also like to thank to all data annotators worked in this project.